OpenStudy (anonymous):

can somone help me find the intersection of these two function, y=x^4 and y=8x

OpenStudy (anonymous):

i did x^4=8x and got

OpenStudy (anonymous):

x(x-2)(x^2+2x+4)

OpenStudy (anonymous):

x^4=8x (0,0) is obivous

OpenStudy (anonymous):

x=2

OpenStudy (anonymous):

right of course

OpenStudy (anonymous):

so i got limits of integration at [0,2], does that qudratic contribute anything

OpenStudy (anonymous):

No

OpenStudy (anonymous):

x^4=8x , x^4 - 8x=0 , x(x^3-8)=0 , x=0 or x^3 - 8 =0 , x=0 or x^3=8 , x=0 or x=2

OpenStudy (anonymous):

get that mathcruncher?

OpenStudy (anonymous):

huh i yeah i got it

OpenStudy (anonymous):

would the 2 be the upper limit of integration and the 0 the lower limit

OpenStudy (anonymous):

Yes, you just need interesection point

OpenStudy (anonymous):

This is a quartic and therefore has 4 roots. by inspection, 0 and 2 are real, the other 2 are complex.

OpenStudy (anonymous):

right, but when i evluated the intergal for the are i got -48/5, the book says i should have gotten +48/5

OpenStudy (anonymous):

x^4 is positive for all x.

OpenStudy (anonymous):

so should i just recongnize that i take the answer as positive

OpenStudy (anonymous):

or is there somthing i am missing

OpenStudy (anonymous):

cause when i graph these function, i see that the area bounded by these functions is in the first qudrant

OpenStudy (anonymous):

I don't know what integral u are referring to. Since x^4 is positive, any area under it must be positive.

OpenStudy (anonymous):

you know the integral formed by those two functions

OpenStudy (anonymous):

evaluated [0,2]

OpenStudy (anonymous):

You are not making any sense, the line 8x is above the curve between 0 and 2.

OpenStudy (anonymous):

right, so i am looking to evluate the area bounded by those two curves

OpenStudy (anonymous):

are u saying u want the area between the curves?

OpenStudy (anonymous):

the area of the region enclosed by the line and curve

OpenStudy (anonymous):

Just calculate the area of the triangle and subtract the integral of 4x^2 between 0 and 2.

OpenStudy (anonymous):

1/2 *h*b

OpenStudy (anonymous):

It's 16 obviously.

OpenStudy (anonymous):

right

OpenStudy (anonymous):

its obviously not 16

OpenStudy (anonymous):

what are you talking about estudier

OpenStudy (anonymous):

i evluated the integral and got -48/5, but the answer is positive 48/5

OpenStudy (anonymous):

\[y=x^4, y=8x\] To find where these two function intersect, set them equal to each other: \[x^4=8x \rightarrow \frac{x^4}{x}=8 \rightarrow x^3=8 \rightarrow \sqrt[3]{x^3}=\sqrt[3]{8} \rightarrow x=2\]

OpenStudy (anonymous):

very good, now take the integral and evluate the are

OpenStudy (anonymous):

can you tell me what you get

OpenStudy (anonymous):

cause my answer is alittle off i guess

OpenStudy (anonymous):

are you trying to find the point where they intersect on the intersection of the areas under both functions?

OpenStudy (anonymous):

The 16 is the area of the triangle, I assumed that's what you meant by 1/2bh. Now u need to subtract the area under x^4 between 0 and 2.

OpenStudy (anonymous):

@nadeem we have already found the intersection points, he want s the area bounded by the line and the curve.

OpenStudy (anonymous):

so, its gonna be 16-35/5

OpenStudy (anonymous):

Which is the triangle formed by the line and the x-axis of area 16 less the area under the curve x^4 between 0 and 2.

OpenStudy (anonymous):

I haven't done the integration, by eye it looks to be about 8 or so.

OpenStudy (anonymous):

okay, but isnt the the integral from 0 to 2, of x^4-8x

OpenStudy (anonymous):

Why do you want to do that?

OpenStudy (anonymous):

find the area

OpenStudy (anonymous):

U already found it, I can see u wrote it up above, 16-7 = 9.

OpenStudy (anonymous):

That's the simplest easiest fastest way to get the area.

OpenStudy (anonymous):

but thats not the answer

OpenStudy (anonymous):

k: \[\int\limits_{0}^{2} 8x-x^4dx = \int\limits_{0}^{2} 8xdx- \int\limits_{0}^{2} x^4dx= \left[ 4x^2 \right]_{0}^{2} - \left[ \frac{x^5}{5} \right]_{0}^{2}= 16-\frac{32}{5}= \frac{48}{5}\]

OpenStudy (anonymous):

thats what im looking for

OpenStudy (anonymous):

how come the function 8x goes before the x^4

OpenStudy (anonymous):

oh itst the upper curve

OpenStudy (anonymous):

because the upper function is 8x and the lover function is x^4, on these kinds of integrals graphing the functions helps tremendously, and it puts everything in perspective as well.

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

sure thing

OpenStudy (anonymous):

16 - int x4 [0,2]^= 16 - x^5/5 -> 16- 32/5 = 48/5 QED.